# Dense $\mathbb{X}$ Retrieval: What Retrieval Granularity Should We Use?

### v2

@@ Abstract
Dense retrieval has become a prominent method to obtain relevant context or world knowledge in open-domain NLP tasks. When we use a learned dense retriever on a retrieval corpus at inference time, an often-overlooked design choice is the retrieval unit in which the corpus is indexed, e.g. document, passage, or sentence. We discover that the retrieval unit choice significantly impacts the performance of both retrieval and downstream tasks. Distinct from the typical approach of using passages or sentences, we introduce a novel retrieval unit, proposition, for dense retrieval. Propositions are defined as atomic expressions within text, each encapsulating a distinct factoid and presented in a concise, self-contained natural language format. We conduct an empirical comparison of different retrieval granularity. Our results reveal that proposition-based retrieval significantly outperforms traditional passage or sentence-based methods in dense retrieval. Moreover, retrieval by proposition also enhances the performance of downstream QA tasks, since the retrieved texts are more condensed with question-relevant information, reducing the need for lengthy input tokens and minimizing the inclusion of extraneous, irrelevant information.
@@ 1 Introduction
Prior to restoration work performed between 1990 and 2001, the tower leaned at an angle of 5.5 degrees, but the tower now leans at about 3.99 degrees. This means the top of the Leaning Tower of Pisa is displaced horizontally 3.9 meters (12 ft 10 in) from the center.
Sentence   Retrieval
@@
Prior to restoration work performed between 1990 and 2001, the tower leaned at an angle of 5.5 degrees, but the tower now leans at about 3.99 degrees.
Proposition   Retrieval
@@
The Leaning Tower of Pisa now leans at about 3.99 degrees.
@@
Figure 1: (Top) An example of three granularities of retrieval units of Wikipedia text when using dense retrieval. (Bottom) We observe that retrieving by propositions yields the best retrieval performance in both passage retrieval task and downstream open-domain QA task, e.g. with Contriever Izacard et al. (2022) or GTR Ni et al. (2022) as the backbone retriever. Highlight indicates the part that contains answer to the question.
@@
Figure 2: We discover that segmenting and indexing a retrieval corpus on the proposition level can be a simple yet effective strategy to increase dense retrievers’ generalization performance at inference time (A, B). We empirically compare the retrieval and downstream open-domain QA tasks performance when dense retrievers work with Wikipedia indexed at the level of 100-word passage, sentence or proposition (C, D).
@@
Dense retrievers are a popular class of techniques for accessing external information sources for knowledge-intensive tasks Karpukhin et al. (2020). Before we use a learned dense retriever to retrieve from a corpus, an imperative design decision we have to make is the retrieval unit – i.e. the granularity at which we segment and index the retrieval corpus for inference. In practice, the choice of retrieval unit, e.g. documents, fixed-length passage chunks or sentences, etc, is usually pre-determined based on how the dense retrieval model is instantiated or trained Lewis et al. (2020); Lee et al. (2021a); Santhanam et al. (2022); Ni et al. (2022).
@@
In this paper, we investigate an overlooked research question with dense retrieval inference – at what retrieval granularity should we segment and index the retrieval corpus? We discover that selecting the proper retrieval granularity at inference time can be a simple yet effective strategy for improving dense retrievers’ retrieval and downstream task performance. We illustrate our intuition with an example of open-domain question-answering (QA) in Figure 1. The example shows retrieved text by the same model at three different granularities. The passage, which represents a coarser retrieval unit with a longer context, is theoretically able to provide more relevant information for the question. However, a passage often includes extraneous details (e.g., restoration period and horizontal displacement in the example of Figure 1) that could potentially distract both the retriever and the language model in downstream tasks Shi et al. (2023); Yu et al. (2023b). On the other hand, sentence-level indexing provides a finer-grained approach but does not entirely address the issue Akkalyoncu Yilmaz et al. (2019); Yang et al. (2020). This is because sentences can still be complex and compounded, and they are often not self-contained, lacking necessary contextual information (e.g., in the example of Figure 1, “the tower” is coreference of “Pisa Tower”) for judging the query-document relevance.
@@
To address these shortcomings of typical retrieval units such as passages or sentences, we propose using proposition as a novel retrieval unit for dense retrieval. Propositions are defined as atomic expressions within text, each encapsulating a distinct factoid and presented in a concise, self-contained natural language format. We show an example proposition in Figure 1. The proposition describes the information regarding the Tower of Pisa’s current leaning angle in a self-contained way and precisely responds to what the question is querying. We provide a more detailed definition and description of proposition in section 2.
@@
To validate the efficacy of using proposition as a retrieval unit for dense retrievers inference, we first process and index an English Wikipedia dump with all documents segmented into propositions, which we refer to as FactoidWiki. Then we conduct experiments on five different open-domain QA datasets and empirically compare the performance of six dual-encoder retrievers when Wikipedia is indexed by passage, sentence, and our proposed proposition. Our evaluation is twofold: we examine both the retrieval performance and the impact on downstream QA tasks. Notably, our findings indicate that proposition-based retrieval outperforms sentence and passage-based methods, especially in terms of generalization, as discussed in section 5. This suggests that propositions, being both compact and rich in context, enable dense retrievers to access precise information while maintaining adequate context. The average improvement over passage-based retrieval of Recall@20 is +10.1 on unsupervised dense retrievers and +2.2 on supervised retrievers. Furthermore, we observe a distinct advantage in downstream QA performance when using proposition-based retrieval, as elaborated in section 6. Given the often limited input token length in language models, propositions inherently provide a higher density of question-relevant information.
Our main contributions in the paper are:
  • We propose using propositions as retrieval units when indexing a retrieval corpus to improve the dense retrieval performance.
@@
  • We introduce FactoidWiki, a processed English Wikipedia dump, where each page is segmented into multiple granularities: 100-word passages, sentences and propositions.
  • We discover that retrieval by proposition outperforms passage or sentence retrieval in terms of generalization for passage retrieval and accuracy for downstream question-answering given the same input token limit.
@@ 2 Proposition as a Retrieval Unit
The goal of our study is to understand how the granularity of a retrieval corpus influences the dense retrieval models’ performance empirically. Aside from commonly-used retrieval units such as 100-word passage Karpukhin et al. (2020) or sentence, we propose using proposition as an alternative retrieval unit choice. Here, propositions represent atomic expressions of meanings in text Min et al. (2023) that are defined by the three principles below.
@@
The use of proposition as a retrieval unit is inspired by a recent line of work Min et al. (2023); Kamoi et al. (2023); Chen et al. (2023a, b), which finds success in representing and evaluating text semantics at the level of propositions. We demonstrate the concept of proposition and how a passage can be split into its set of propositions by an example on the left side of Figure 2. The passage contains three propositions, each of which corresponds to a distinct factoid about the Leaning Tower of Pisa: the angle before the restoration, the current angle, and the horizontal displacement. Within each proposition, necessary context from the passage is incorporated so that the meaning of the proposition can be interpreted independently of the original text, e.g. the reference of the tower is resolved into its full mention, the Leaning Tower of Pisa, in the first proposition. We expect each proposition to describe exactly one contextualized atomic fact, and so our intuition is that propositions would suitably work as a retrieval unit for information-seeking questions.
@@ 3 FactoidWiki: Proposition-Level Index and Retrieval for Wikipedia
We empirically compare the use of 100-word passages, sentences, and propositions as retrieval units on Wikipedia, a commonly-used retrieval source for knowledge-intensive NLP tasks Petroni et al. (2021). To allow for a fair comparison across granularities, we process an English Wikipedia dump from 2021-10-13, as used by Bohnet et al. (2022). We segment each document text into three different granularities: 100-word passages, sentences, and propositions. We include the details on passage- and sentence-level segmentation of the corpus in Appendix A.
@@
To segment the Wikipedia pages into propositions, we finetune a text generation model, which we refer to as the Propositionizer. The Propositionizer takes a passage as input and generates the list of propositions within the passage. Following Chen et al. (2023b), we train the Propositionizer with a two-step distillation process. We first prompt GPT-4 OpenAI (2023) with an instruction containing the proposition definition and 1-shot demonstrations. We include the details of the prompt in Figure 8. We start with a set of 42k passages and use GPT-4 to generate the seed set of paragraph-to-propositions pairs. Next, we use the seed set to finetune a Flan-T5-large model Chung et al. (2022).
We refer to the processed corpus as FactoidWiki. The resulting statistics of FactoidWiki are shown in Table 1.
@@
Passage 41,393,528 58.5
Sentence 114,219,127 21.0
Proposition 256,885,003 11.2
Table 1: Statistics of text units in the English Wikipedia dump from 2021-10-13.
@@ 4.1 Open-Domain QA Datasets
We evaluate on five different open-domain QA datasets with Wikipedia as the retrieval source: Natural Questions (NQ) (Kwiatkowski et al., 2019), TriviaQA (TQA) (Joshi et al., 2017), Web Questions (WebQ)  (Berant et al., 2013), SQuAD (Rajpurkar et al., 2016), and Entity Questions (EQ) (Sciavolino et al., 2021).
@@ 4.2 Dense Retrieval Models
We compare the performance of the six following supervised or unsupervised dense retriever models. Here, supervised models refer to ones that have used human-labeled query-passage pairs as supervision during training, and vice versa.
  • SimCSE (Gao et al., 2021) is a BERT-base Devlin et al. (2019) encoder trained on unlabeled sentence randomly sampled from Wikipedia.
@@
  • DPR (Karpukhin et al., 2020) is a dual-encoder BERT-base model finetuned on five open-domain QA datasets, which includes four of the datasets (NQ, TQA, WebQ and SQuAD) in our evaluation.
  • ANCE (Xiong et al., 2020) used the same setting from DPR and trained the model with Approximate nearest neighbor Negative Contrastive Estimation (ANCE), a hard negatives-based training approach.
  • TAS-B (Hofstätter et al., 2021) is a dual-encoder BERT-base model distilled from a teacher model with cross-attention trained on MS MARCO (Nguyen et al., 2016).
  • GTR (Hofstätter et al., 2021) is a T5-base encoder Raffel et al. (2020) pretrained on unlabeled pairs of online forum QA data, and fine-tuned on MS MARCO and Natural Question.
@@ 4.3 Passage Retrieval Evaluation
We evaluate retrieval performance at the passage level when the corpus is indexed at the passage, sentence, or proposition level respectively. For sentence and proposition level retrieval, we follow the setting introduced in Lee et al. (2021b), where the score of the passage is based on the maximum similarity score between the query and all sentences or propositions in a passage. In practice, we first retrieve a slightly larger number of text units, map each unit to the source passage, and return the top-$k$ unique passages. We use Recall@$k$ as our evaluation metric, which is defined as the percentage of questions for which the correct answer is found within the top-$k$ retrieved passages.
4.4 Downstream QA Evaluation
To understand the implications of using different retrieval units on the downstream open-domain QA tasks, we evaluate the use of retrieval models in retrieve-then-read setup Izacard and Grave (2021). With the retrieve-then-read setting, a retrieval model first retrieves $k$ text units given the query. The $k$ retrieved text units are then used as input along with the query to a reader model to derive the final answer. Typically, the choice of $k$ is subject to the reader model’s maximum input length constraint, or the limit of compute budget, which scales with the number of input tokens.
For this reason, we follow an evaluation setup where the maximum number of retrieved words is capped at $l=100$ or $500$, i.e. only the top $l$ words from passage, sentence, or proposition level retrieval are feed into the reader model as input. We evaluate the percentage of questions for which the predicted answer exactly matches (EM) the ground truth. We denote our metric as EM @ $l$. For our evaluation, we use T5-large size UnifiedQA-v2 as the reader model  (Khashabi et al., 2022).
5 Results: Passage Retrieval
In this section, we report and discuss the retrieval tasks performance. Our results show that despite none of the models being trained with proposition-level data, all the retrieval models demonstrated on-par or superior performance when the corpus is indexed at the proposition level.
5.1 Passage Retrieval Performance
@@
ANCE Passage 70.7 81.4 73.9 81.4 65.7 77.2 43.3 58.6 57.0 69.1 62.1 73.5
Sentence 70.3 81.6 73.9 81.5 65.2 77.4 45.8 60.7 61.4 72.8 63.3 74.8
Proposition 69.9 81.1 72.8 80.6 65.1 77.1 46.2 61.9 66.7 76.6 64.1 75.5
TAS-B Passage 64.2 77.9 70.4 79.3 65.1 77.0 54.3 69.2 72.2 81.3 65.2 76.9
Sentence 64.0 78.4 71.4 80.2 63.9 76.7 58.9 72.3 72.7 82.0 66.2 77.9
Proposition 63.8 78.6 71.4 80.0 63.8 76.8 59.8 73.4 75.1 83.3 66.8 78.4
@@
Figure 3: Document retrieval recall vs. the popularity of the target entity in each question from the EntityQuestions dataset. The popularity of each entity (i.e. smaller value $\Rightarrow$ less common entities, and vice versa) is estimated by the occurrence of the entity in its top-1000 passage retrieved by BM25. On queries with less common entities, we observe that retrieving by proposition shows a larger advantage over retrieval by proposition.

### v3

# EMNLP 2024 Main Conference Dense $\mathbb{X}$ Retrieval: What Retrieval Granularity Should We Use?

@@ 1 Introduction
Passage   Retrieval Prior to restoration work performed between 1990 and 2001, the tower leaned at an angle of 5.5 degrees, but the tower now leans at about 3.99 degrees. This means the top of the Leaning Tower of Pisa is displaced horizontally 3.9 meters (12 ft 10 in) from the center.
@@
Sentence   Retrieval Prior to restoration work performed between 1990 and 2001, the tower leaned at an angle of 5.5 degrees, but the tower now leans at about 3.99 degrees.
@@
Proposition   Retrieval The Leaning Tower of Pisa now leans at about 3.99 degrees.
@@
Figure 1: (Top) An example of three granularities of retrieval units of Wikipedia text when using dense retrieval. (Bottom) We observe that retrieving by propositions yields the best retrieval performance in both passage retrieval task and downstream open-domain QA task, e.g. with Contriever Izacard et al. (2022) or GTR Ni et al. (2022) as the backbone retriever. Highlight indicates the part that contains the answer to the question.
@@
Figure 2: We discover that segmenting and indexing a retrieval corpus on the proposition level can be a simple yet effective strategy to increase dense retrievers’ generalization performance at inference time (A, B). We empirically compare the retrieval and downstream open-domain QA task performance when dense retrievers work with Wikipedia indexed at the level of 100-word passages, sentences, or propositions (C, D).
@@
Dense retrievers are a popular class of techniques for accessing external information sources for open-domain NLP tasks Karpukhin et al. (2020). Before we use a learned dense retriever to retrieve from a corpus, an imperative design decision we have to make is the retrieval unit – i.e. the granularity at which we segment and index the retrieval corpus for inference. In practice, the choice of retrieval units, e.g. documents, fixed-length passage chunks or sentences, etc, is usually pre-determined based on how the dense retrieval model is instantiated or trained Lewis et al. (2020); Lee et al. (2021a); Santhanam et al. (2022); Ni et al. (2022).
In this paper, we investigate an overlooked research question with dense retrieval inference – at what retrieval granularity should we segment and index the retrieval corpus? We aim to investigate this question in two aspects.
$\bullet$ First, we examine how the granularity of the index affects passage retrieval performance.
$\bullet$ Second, we investigate whether fine-grained units can replace passages in downstream QA tasks.
@@
Based on our empirical experiments, we discover that selecting the proper retrieval granularity at inference time can be a simple yet effective strategy for improving dense retrievers’ retrieval and downstream QA performance. We illustrate our intuition with an example of open-domain QA in Figure 1. The example shows retrieved text by the same retriever at three different granularities. The passage, which represents a coarser retrieval unit with a longer context, is theoretically able to provide more relevant information for the question. However, a passage often includes extraneous details (e.g., restoration period and horizontal displacement in the example of Figure 1) that could potentially distract both the retriever and the language model in downstream tasks Shi et al. (2023); Yu et al. (2023b). On the other hand, sentence-level indexing provides a finer-grained approach but does not entirely address the issue Akkalyoncu Yilmaz et al. (2019); Yang et al. (2020). This is because sentences can still be complex and compounded, and they are often not self-contained, lacking necessary contextual information (e.g., in the example of Figure 1, “the tower” is the coreference of “Pisa Tower”) for judging the query-document relevance.
@@
To address these shortcomings of typical retrieval units such as passages or sentences, we propose using proposition as a novel retrieval unit for dense retrieval. Propositions are defined as atomic expressions within text, where each encapsulates a distinct factoid and is presented in a concise, self-contained natural language format. We show an example proposition in Figure 1. The proposition describes the information regarding the Tower of Pisa’s current leaning angle in a self-contained way and precisely responds to what the question is querying. We provide a more detailed definition and description of proposition in section 2. To validate the efficacy of using proposition as a retrieval unit for dense retrievers inference, we first process and index an English Wikipedia dump with all documents segmented into propositions, which we refer to as FactoidWiki.
@@
We conduct experiments on five different open-domain QA datasets and empirically compare the performance of four dual-encoder retrievers when Wikipedia is indexed by passages, sentences, and our proposed propositions. Notably, our findings indicate that proposition-based retrieval outperforms sentence and passage-based retrieval, especially in terms of generalization, as discussed in section 5. This suggests that propositions, being both compact and rich in context, enable dense retrievers to access precise information while maintaining adequate context. The average improvement over passage-based retrieval of Recall@20 is +10.1 on unsupervised dense retrievers and +2.7 on supervised retrievers, even though these retrievers were directly trained on passage-level retrieval. Furthermore, we observe a distinct advantage of proposition-based retrieval in downstream QA performance when using retrieval-augmented language models, as elaborated in  section 6. Retrieval by finer-grained units inherently provides a higher density of question-relevant information. This finding implies using finer-grained units in the prompts achieves the same performance with a shorter input length, and hence, a faster inference time.
Our main contributions are:
  • We provide a systemic study on how retrieval granularity impacts retrieval and downstream task performance. We observe that the retrieval units have a significant impact on performance.
@@
  • We introduce FactoidWiki, a processed English Wikipedia dump, where each page is segmented into multiple granularities: passages, sentences, and our proposed propositions.
  • We propose retrieval by proposition as an alternative strategy, which achieves better retrieval and QA accuracy and generalization performance (with unsupervised retriever), compared to passage or sentence as retrieval unit.
@@ 2 Proposition as a Retrieval Unit
The goal of our study is to understand how the granularity of a retrieval corpus influences the dense retrieval models’ performance empirically. Aside from commonly-used retrieval units such as 100-word passages Karpukhin et al. (2020) or sentences, we propose using proposition as an alternative retrieval unit choice. Here, propositions represent atomic expressions of meanings in text Min et al. (2023) with three defining principles below.
@@
The use of proposition as a retrieval unit is inspired by a recent line of work Min et al. (2023); Kamoi et al. (2023); Chen et al. (2023a, b), which finds success in representing and evaluating text semantics at the level of propositions. We demonstrate the concept of proposition and how a passage can be split into a set of propositions by an example on the left side of Figure 2. The passage contains three propositions, each of which corresponds to a distinct factoid about the Leaning Tower of Pisa: the angle before the restoration, the current angle, and the horizontal displacement.
Within each proposition, necessary context from the passage is incorporated so that the meaning of the proposition can be interpreted independently of the original text, e.g. the reference of the tower is resolved into its full mention, the Leaning Tower of Pisa, in the first proposition. We expect each proposition to describe exactly one atomic fact, and so our intuition is that propositions would suitably work as a retrieval unit for information-seeking questions.
@@ 3 FactoidWiki: Proposition-Level Index and Retrieval for Wikipedia
We empirically compare passages, sentences, and propositions as retrieval units on Wikipedia, a commonly-used retrieval source for knowledge-intensive NLP tasks Petroni et al. (2021). To allow a fair comparison across granularities, we process an English Wikipedia dump from 2021-10-13, as used by Bohnet et al. (2022). We segment each document text into three different granularities: passages, sentences, and propositions. We include the details on passage- and sentence-level segmentation of the corpus in Appendix A.
@@
To segment the Wikipedia pages into propositions, we finetune a text generation model, which we refer to as the Propositionizer. The Propositionizer takes a passage as input and generates the list of propositions within the passage.
Following Chen et al. (2023b), we train the Propositionizer with a two-step distillation process. We first prompt GPT-4 OpenAI (2023) with an instruction containing the proposition definition and 1-shot demonstration. We include the details of the prompt in Figure 8. We start with a set of 42k passages and use GPT-4 to generate the seed set of paragraph-to-proposition pairs. Next, we use the seed set to finetune a Flan-T5-large model Chung et al. (2022). We refer to the processed corpus as FactoidWiki. The statistics of FactoidWiki are shown in Table 1.
@@
Passages   41,393,528 58.5
Sentences 114,219,127 21.0
Propositions 256,885,003 11.2
Table 1: Statistics of text units in the English Wikipedia.
Quality Analysis.
We conduct a manual error analysis to understand the quality of propositions generated by GPT-4 and the Propositionizer. While there does not exist a fixed standard on deciding a ground truth set of propositions for a passage, we estimate the frequency of error cases where (1) a proposition is not fully supported by the passage, (2) a proposition can be further split into separate propositions, and (3) propositions are not self-contained, respectively (Table 2). On a random sample of 50 passages, we observe that almost all propositions generated by both models are faithful, while a small portion of the propositions are not stand-alone.
GPT-4 Propositionizer
Not Faithful 0.7% (3/408) 1.3% (6/445)
Not Minimal 2.9% (12/408) 2.0% (9/445)
Not Stand-alone 4.9% (20/408) 3.1% (14/445)
Table 2: Frequency of errors occurred in the generated propositions. Most generated propositions are faithful, while a small portion of them are not stand-alone.
@@ 4.1 Open-Domain QA Datasets
We experiment on five different open-domain QA datasets with Wikipedia as the retrieval source: Natural Questions  (NQ, Kwiatkowski et al., 2019), TriviaQA  (TQA, Joshi et al., 2017), Web Questions  (WebQ, Berant et al., 2013), SQuAD (Rajpurkar et al., 2016), and Entity Questions (EQ, Sciavolino et al., 2021).
@@ 4.2 Dense Retrieval Models
We compare the performance of the four following supervised or unsupervised dense retriever models. Here, supervised models refer to ones that have used human-labeled query-passage pairs as supervision during training, and vice versa.
  • SimCSE (Gao et al., 2021) is a BERT-base Devlin et al. (2019) encoder trained on unlabeled sentences sampled randomly from Wikipedia. SimCSE can be transferred to use as an unsupervised retriever (Chen et al., 2023b).
@@
  • DPR (Karpukhin et al., 2020) is a dual-encoder BERT-base model fine-tuned on passage retrieval tasks directly using the question-passage pair labels from NQ, TQA, WebQ and SQuAD.
  • GTR (Ni et al., 2022) is a T5-base encoder Raffel et al. (2020) pretrained on online forum QA data, and fine-tuned with question-passage pair labels on MS MARCO (Nguyen et al., 2016) and NQ datasets.
@@ 4.3 Passage Retrieval Evaluation
We evaluate the retrieval performance at the passage level when the corpus is indexed at the passage, sentence, or proposition level respectively. For sentence and proposition level retrieval, we follow the setting introduced in Lee et al. (2021b), where the score of the passage is based on the maximum similarity score between the query and all sentences or propositions in a passage. In practice, we first retrieve a slightly larger number of text units, then map each unit to the source passage, and eventually return the top-$k$ unique passages. We use Passage Recall@$k$ as our evaluation metric, which is defined as the percentage of questions for which the correct answer is found within the top-$k$ retrieved passages.
To further understand how different retrieved passages affect the downstream QA. We use Fusion-in-Decoder (FiD, Izacard and Grave, 2021) model to extract answers from retrieved passages. We use a T5-large sized FiD model trained on NQ dataset in our experiments. The exact match (EM) score computes the percentage of questions for which the predicted answer exactly matches the ground truth.
@@
4.4 Open-domain QA Evaluation on Retrieval-Augmented Language Models
Another aspect of the choice of granularity lies in what units should be used in the prompt for retrieval-augmented language models. For large language models, retrieval-augmented generation is achieved by prepending retrieved units to user instruction and taking them as the input for language models. We aim to understand the implications of using retrieved units of different granularity within the same computational budget at inference time. To fairly compare using different granularity in the prompts under the same computation budget, we set a token length limit for retrieved units.
For this reason, we follow an evaluation setup where the maximum number of retrieved tokens is capped at $l=100$ or $500$, i.e. only the top $l$ tokens from passage, sentence, or proposition level retrieval are fed into the language model as input. We evaluate the percentage of questions for which the predicted answer exactly matches (EM) the ground truth. We denote our metric as EM @ $l$ tokens. We use LLaMA-2-7B (Touvron et al., 2023) in our evaluation. To ensure the model’s output aligns with the format of each dataset, we employ in-context learning, incorporating four-shot demonstrations as illustrated in Figure 9.
@@ 4.4 Open-domain QA Evaluation on Retrieval-Augmented Language Models
Figure 3: Document retrieval recall vs. the frequency of the target entity in each question from the Entity Questions dataset. The frequency of each entity (i.e. smaller value $\Rightarrow$ less common entities, and vice versa) is estimated by the frequency of the entity in its top-1000 passage retrieved by BM25. On queries with less common entities, we observe that retrieving by proposition shows a larger advantage over retrieval by proposition.
Retriever Granularity NQ TQA WebQ SQuAD EQ Avg.
top-5 top-20 top-5 top-20 top-5 top-20 top-5 top-20 top-5 top-20 top-5 top-20
Unsupervised Dense Retrievers
SimCSE Passage 16.6 23.6 32.3 40.8 15.5 19.1 14.6 20.7 16.1 20.3 19.0 24.9
Sentence 20.7 28.1 36.0 44.5 18.5 21.9 19.6 25.8 19.9 25.1 23.0 29.1
Proposition 24.5 33.1 37.5 46.2 19.7 23.0 21.4 27.6 26.8 32.0 26.0 32.4
Contriever Passage 23.2 35.1 40.8 50.8 16.3 22.1 23.9 32.7 20.2 27.9 24.9 33.7
Sentence 26.0 36.8 43.4 52.9 18.4 23.9 26.7 34.7 23.7 30.3 27.6 35.7
Proposition 28.9 39.2 47.2 55.6 19.5 25.2 30.8 37.6 28.8 35.8 31.1 38.7
Supervised Dense Retrievers
DPR Passage 41.1 45.6 50.6 57.0 23.7 25.5 18.8 25.4 25.3 29.7 31.9 36.6
Sentence 40.3 45.6 51.7 57.6 24.0 26.9 21.1 27.4 28.6 32.9 33.1 38.1
Proposition 39.7 45.2 51.0 56.8 24.3 27.5 22.2 28.3 32.0 36.0 33.9 38.8
GTR Passage 39.8 46.1 49.7 55.9 23.0 25.9 29.9 35.1 37.8 39.6 36.0 40.5
Sentence 39.4 45.9 51.7 58.0 23.2 26.1 35.7 39.1 38.0 39.9 37.6 41.8
Proposition 40.0 46.9 52.5 58.4 24.2 26.5 37.8 40.4 39.2 41.0 38.7 42.6
Table 4: Open-domain QA performance (Exact Match) using Fusion-in-Decoder model (Izacard and Grave, 2021) to extract answer from top-5 and top-20 passages retrieved on the index of passages, sentences, and propositions.
5 How Does Granularity Influence Passage Retrieval?
In this section, we report and discuss how indexing the corpus at various granularity influences the passage retrieval performance. Surprisingly, despite all of the dense retrieval models being trained on only passage-level documents, all the models demonstrate on-par or superior performance when the corpus is indexed at the proposition level. Our results suggest that indexing the corpus at the finer-grained units improves the cross-task generalization on passage retrieval.
5.1 Passage Retrieval Performance